# OATS: Outlier-Aware Pruning through Sparse and Low Rank Decomposition

### v1

@@ A.2 MMLU Subjects
A.3 Attention Rollout: Details
@@
A.4 Pruning Implementation and Hyperparameters
@@ A.4.1 SparseGPT Hyperparameters
A.4.2 DSNoT Hyperparameters
We run experiments utilizing DSNoT where the initial masks are generated by SparseGPT and Wanda. All DSNoT experiments were run with $50$ iterations and an update threshold of $0.1$. Depicted in Table 10 below show the results distinguishing between the two initial methods that were utilized.
Model Compression Rate Method MMLU ($\uparrow$) Avg. Zero-Shot($\uparrow$) Perplexity($\downarrow$) Phi-3 Mini 30% DSNoT w/ SparseGPT 67.01 70.81 10.55 DSNoT w/ Wanda 68.02 71.20 10.51 40% DSNoT w/ SparseGPT 62.94 68.86 12.29 DSNoT w/ Wanda 63.57 69.08 12.17 50% DSNoT w/ SparseGPT 53.99 64.74 16.71 DSNoT w/ Wanda 54.28 65.33 16.68 Phi-3 Medium 30% DSNoT w/ SparseGPT 74.89 73.82 7.11 DSNoT w/ Wanda 75.13 74.03 7.11 40% DSNoT w/ SparseGPT 73.15 72.54 8.24 DSNoT w/ Wanda 73.20 72.90 8.27 50% DSNoT w/ SparseGPT 68.65 71.12 9.96 DSNoT w/ Wanda 68.12 71.10 10.02 Llama-3 8B 30% DSNoT w/ SparseGPT 62.99 68.98 9.37 DSNoT w/ Wanda 63.72 68.64 9.36 40% DSNoT w/ SparseGPT 58.97 66.28 9.60 DSNoT w/ Wanda 59.99 66.65 9.68 50% DSNoT w/ SparseGPT 49.15 62.74 12.41 DSNoT w/ Wanda 49.20 62.35 12.42
@@ A.4.2 DSNoT Hyperparameters
Table 11 below depicts the analogous results but for our vision transformer experiments:
Model Compression Rate Method Accuracy (%) ViT-Base 30% DSNoT w/ SparseGPT 80.01 DSNoT w/ Wanda 80.16 40% DSNoT w/ SparseGPT 79.12 DSNoT w/ Wanda 79.46 50% DSNoT w/ SparseGPT 75.83 DSNoT w/ Wanda 76.90 DinoV2-Giant 30% DSNoT w/ SparseGPT 86.46 DSNoT w/ Wanda 86.45 40% DSNoT w/ SparseGPT 86.37 DSNoT w/ Wanda 86.30 50% DSNoT w/ SparseGPT 85.87 DSNoT w/ Wanda 85.93

### v2

@@ A.9 MMLU Subjects
A.10 Attention Rollout: Details
@@
A.11 Zero-Shot Task-Specific Performance
Table 17, below, shows the task-specific performance for the zero-shot evaluation results presented in Section 3.2 and Appendix A.13.
Model Compression Method PIQA HellaSwag WinoGrande OpenBookQA RTE BoolQ ARC-e ARC-c Phi-3 Mini 0% Dense 81.23 77.50 73.56 46.80 75.81 85.32 78.45 57.25 30% SparseGPT 78.94 76.94 69.85 49.60 73.29 84.13 76.39 55.89 Wanda 79.65 76.27 71.59 48.00 73.65 83.70 77.23 55.20 DSNoT w/ SparseGPT 80.09 75.61 72.22 47.40 74.37 84.22 77.86 54.69 DSNoT w/ Wanda 80.41 75.52 72.06 47.60 74.37 84.53 79.55 55.55 OATS 80.03 77.07 72.61 47.60 74.37 84.92 77.44 57.76 40% SparseGPT 78.35 75.07 68.59 47.00 72.20 83.67 75.29 53.24 Wanda 78.35 73.87 69.30 45.40 71.84 83.18 76.52 51.96 DSNoT w/ SparseGPT 78.56 72.99 70.64 46.20 70.40 82.72 76.52 52.82 DSNoT w/ Wanda 79.33 73.06 70.88 44.00 70.76 83.79 77.40 53.41 OATS 79.38 75.86 70.01 46.60 72.56 83.98 76.85 55.12 50% SparseGPT 77.20 70.63 66.46 45.20 70.76 83.06 70.58 47.01 Wanda 76.33 67.70 66.38 41.80 66.43 81.83 72.43 47.35 DSNoT w/ SparseGPT 76.28 67.16 65.90 42.20 63.90 81.56 72.90 48.04 DSNoT w/ Wanda 75.52 66.54 67.64 43.00 65.34 82.54 73.48 48.55 OATS 77.26 71.64 69.53 44.80 73.65 81.28 77.10 52.05 Phi-3 Medium 0% Dense 81.66 82.83 75.85 50.00 77.62 88.17 78.41 59.64 30% SparseGPT 81.39 82.02 75.77 50.80 77.26 87.80 80.05 61.18 Wanda 81.39 80.88 76.01 49.40 76.90 87.74 79.59 60.49 DSNoT w/ SparseGPT 81.94 80.76 76.95 48.40 75.81 87.65 79.25 59.81 DSNoT w/ Wanda 81.66 81.03 77.27 49.20 76.53 87.80 78.96 59.81 OATS 81.07 82.09 74.43 51.20 78.34 88.38 78.16 58.70 40% SparseGPT 80.41 80.70 75.53 51.20 77.26 88.32 81.23 60.58 Wanda 79.87 78.15 75.45 48.60 77.26 87.71 78.11 58.96 DSNoT w/ SparseGPT 79.82 78.07 75.37 47.00 76.53 87.98 77.31 58.19 DSNoT w/ Wanda 80.30 78.11 74.66 47.80 77.26 88.04 78.11 58.87 OATS 81.39 81.72 75.06 51.00 77.62 87.65 80.39 60.84 50% SparseGPT 79.71 78.27 73.64 50.40 75.45 87.09 82.03 59.39 Wanda 78.29 74.07 74.03 45.00 75.81 85.72 77.44 57.34 DSNoT w/ SparseGPT 79.27 74.30 74.59 44.40 76.90 85.26 77.69 56.57 DSNoT w/ Wanda 78.56 73.81 75.14 43.60 75.81 86.33 77.53 58.02 OATS 81.07 79.18 76.09 50.20 74.73 87.77 80.05 58.02 Llama-3 8B 0% Dense 80.74 79.16 73.40 45.0 0 67.87 80.98 77.69 53.50 30% SparseGPT 80.36 78.58 73.24 44.40 66.79 81.38 76.81 51.11 Wanda 79.98 78.00 73.64 44.40 64.26 81.62 76.18 50.94 DSNoT w/ SparseGPT 80.20 78.12 73.80 44.40 65.70 82.20 75.72 51.71 DSNoT w/ Wanda 79.82 77.99 73.09 44.80 63.18 81.80 77.06 51.37 OATS 80.03 78.75 73.64 45.20 66.06 81.13 76.94 52.99 40% SparseGPT 79.16 76.74 73.32 41.80 64.26 81.31 74.71 49.32 Wanda 78.73 75.90 72.22 44.40 63.18 80.46 72.31 49.15 DSNoT w/ SparseGPT 78.29 75.92 73.32 42.60 58.48 80.86 73.11 47.70 DSNoT w/ Wanda 78.51 75.52 73.24 43.80 61.73 80.70 72.01 47.70 OATS 79.71 77.18 74.19 43.80 67.51 82.39 74.92 49.74 50% SparseGPT 77.58 73.12 72.85 40.80 59.21 79.30 69.28 45.14 Wanda 77.53 69.34 70.24 40.00 61.73 76.57 66.96 43.77 DSNoT w/ SparseGPT 76.88 69.45 69.30 39.60 59.21 77.25 67.93 43.32 DSNoT w/ Wanda 77.09 68.57 69.77 38.60 57.76 76.27 67.34 43.43 OATS 77.75 73.17 71.74 41.00 64.98 79.66 72.35 45.05 Llama-3 70B 0% Dense 84.33 84.89 80.35 48.60 68.23 85.26 86.03 64.51 30% SparseGPT 84.66 84.63 80.35 48.00 69.31 85.26 85.02 63.31 Wanda 84.39 83.97 80.58 48.40 70.04 85.29 85.06 63.82 DSNoT w/ SparseGPT 84.06 84.49 80.11 48.20 69.68 85.57 85.10 63.82 DSNoT w/ Wanda 84.55 84.48 81.22 47.80 71.12 85.93 85.06 64.16 OATS 84.28 84.40 80.66 48.40 69.31 85.32 85.90 63.65 40% SparseGPT 83.62 83.77 80.03 47.80 69.68 85.69 84.47 61.95 Wanda 83.57 83.03 78.93 47.40 68.23 85.05 84.34 62.29 DSNoT w/ SparseGPT 82.37 83.21 78.85 46.20 66.43 85.20 83.75 60.07 DSNoT w/ Wanda 83.79 83.35 79.72 46.80 67.87 85.57 84.89 62.37 OATS 84.44 83.69 80.11 48.60 70.40 84.56 84.55 62.71 50% SparseGPT 83.13 81.68 79.32 46.20 71.12 85.17 81.27 57.51 Wanda 83.08 81.12 78.22 48.00 69.31 84.22 81.61 57.25 DSNoT w/ SparseGPT 81.34 80.68 77.82 45.60 70.04 84.62 80.98 55.12 DSNoT w/ Wanda 85.24 81.64 78.45 46.80 69.31 85.23 81.52 57.76 OATS 83.41 82.16 79.01 47.40 68.59 85.47 82.11 58.28
Table 17: Task-Specific Zero-Shot Results
A.12 Wikitext-2 Perplexity
We utilized LM Harness (Gao et al., 2024) to measure the Wikitext-2 perplexities for the Phi-3 Mini, Medium, and Llama-3 8B models in Table 4. However, prior works (Sun et al., 2024b; Zhang et al., 2024b) have also opted to measure Wikitext-2 perplexities utilizing a different normalization and dataset segmentation. We report the Wikitext-2 perplexities utilizing the latter  [^2] We use the implementation sourced from the code base of Sun et al. (2024b): https://github.com/locuslab/wanda/blob/main/lib/eval.py approach in Table 18 below:
Compression Method Phi-3 Llama-3
Mini ($\mathbf{3.8}$B) Medium ($\mathbf{14}$B) 8B
0% Dense 5.64 4.02 5.54
30% SparseGPT 6.32 4.56 6.04
Wanda 6.19 4.46 6.03
DSNoT 6.11 4.40 5.99
OATS 5.96 4.30 5.87
40% SparseGPT 7.05 5.02 6.75
Wanda 6.99 4.97 6.69
DSNoT 6.81 4.89 6.67
OATS 6.51 4.64 6.39
50% SparseGPT 8.55 5.65 8.30
Wanda 8.84 5.73 8.79
DSNoT 8.58 5.68 8.84
OATS 7.98 5.28 7.77
Table 18: Comparison of perplexity (lower is better) utilizing a different normalization and dataset segmentation.
A.13 Pruning Implementation and Hyperparameters
@@ A.13.1 SparseGPT Hyperparameters
A.13.2 DSNoT Hyperparameters
We run experiments utilizing DSNoT where the initial masks are generated by SparseGPT and Wanda. All DSNoT experiments were run with $50$ iterations and an update threshold of $0.1$. Table 19, below, shows the results distinguishing between the two initial methods that were utilized.
Model Compression Method MMLU ($\uparrow$) Zero-Shot($\uparrow$) Perplexity($\downarrow$)
Phi-3 Mini 30% DSNoT w/ SparseGPT 67.01 70.81 10.55
DSNoT w/ Wanda 68.02 71.20 10.51
40% DSNoT w/ SparseGPT 62.94 68.86 12.29
DSNoT w/ Wanda 63.57 69.08 12.17
50% DSNoT w/ SparseGPT 53.99 64.74 16.71
DSNoT w/ Wanda 54.28 65.33 16.68
Phi-3 Medium 30% DSNoT w/ SparseGPT 74.89 73.82 7.11
DSNoT w/ Wanda 75.13 74.03 7.11
40% DSNoT w/ SparseGPT 73.15 72.54 8.24
DSNoT w/ Wanda 73.20 72.90 8.27
50% DSNoT w/ SparseGPT 68.65 71.12 9.96
DSNoT w/ Wanda 68.12 71.10 10.02
Llama-3 8B 30% DSNoT w/ SparseGPT 62.99 68.98 9.37
DSNoT w/ Wanda 63.72 68.64 9.36
40% DSNoT w/ SparseGPT 58.97 66.28 9.60
DSNoT w/ Wanda 59.99 66.65 9.68
50% DSNoT w/ SparseGPT 49.15 62.74 12.41
DSNoT w/ Wanda 49.20 62.35 12.42
Llama-3 70B 30% DSNoT w/ SparseGPT 78.76 75.13 3.28
DSNoT w/ Wanda 79.00 75.54 3.27
40% DSNoT w/ SparseGPT 76.39 73.26 4.16
DSNoT w/ Wanda 77.70 74.29 4.10
50% DSNoT w/ SparseGPT 72.18 72.02 5.87
DSNoT w/ Wanda 72.76 72.91 5.58
@@ A.13.2 DSNoT Hyperparameters
Table 20, below, shows the analogous results but for our vision transformer experiments:
Model Compression Method Accuracy (%)
ViT-Base 30% DSNoT w/ SparseGPT 80.01
DSNoT w/ Wanda 80.16
40% DSNoT w/ SparseGPT 79.12
DSNoT w/ Wanda 79.46
50% DSNoT w/ SparseGPT 75.83
DSNoT w/ Wanda 76.90
DinoV2-Giant 30% DSNoT w/ SparseGPT 86.46
DSNoT w/ Wanda 86.45
40% DSNoT w/ SparseGPT 86.37
DSNoT w/ Wanda 86.30
50% DSNoT w/ SparseGPT 85.87
DSNoT w/ Wanda 85.93